# A Conflicts-free, Speed-lossless KAN-based Reinforcement Learning Decision System for Interactive Driving in Roundabouts

### v1

@@ Abstract
Safety and efficiency are crucial for autonomous driving in roundabouts, especially in the context of mixed traffic where autonomous vehicles (AVs) and human-driven vehicles coexist. This paper introduces a learning-based algorithm tailored to foster safe and efficient driving behaviors across varying levels of traffic flows in roundabouts. The proposed algorithm employs a deep Q-learning network to effectively learn safe and efficient driving strategies in complex multi-vehicle roundabouts. Additionally, a KAN (Kolmogorov-Arnold network) enhances the AVs’ ability to learn their surroundings robustly and precisely. An action inspector is integrated to replace dangerous actions to avoid collisions when the AV interacts with the environment, and a route planner is proposed to enhance the driving efficiency and safety of the AVs. Moreover, a model predictive control is adopted to ensure stability and precision of the driving actions. The results show that our proposed system consistently achieves safe and efficient driving whilst maintaining a stable training process, as evidenced by the smooth convergence of the reward function and the low variance in the training curves across various traffic flows. Compared to state-of-the-art benchmarks, the proposed algorithm achieves a lower number of collisions and reduced travel time to destination.
@@ I Introduction
As urban roadways evolve, roundabouts have significantly increased the vehicle distribution and road capacity [1]. Compared to other crucial traffic scenarios, roundabouts have been verified to offer fewer conflicts [2]. However, the safety issue becomes more pronounced in high-traffic roundabouts, which has a higher potential for crashes [3]. The design of roundabouts varies for cities of different scales [4]. Typically, a roundabout is a type of circular intersection created to enable a smooth flow of traffic, characterized by a central island that is not meant for vehicular access [5]. In standard roundabouts, the direction of travel is set—either clockwise or counterclockwise—simplifying the interaction complexities. The essential driving maneuvers in a roundabout include entering, lane selection, circulating, lane-changing, and exiting. Both merging into and exiting from the roundabout require interaction between the host vehicles and surrounding human-driven vehicles (HDVs), with host vehicles making self-optimized decisions by interpreting the intentions of other drivers. Choosing the correct lane allows host vehicles to maneuver with greater flexibility and efficiency. The lane-changing process obliges host vehicles to carefully monitor the state of surrounding HDVs to avoid possible collisions. Precise control is essential to adhere to the generated trajectories throughout the driving process. Thus, during navigation in roundabouts, host vehicles must have the ability to select the appropriate lane, monitor nearby entities and the environment, avoid collisions, and precisely control their movement.
Autonomous vehicles (AVs) can significantly reduce safety incidents that stem from human errors such as fatigue, distraction, and delayed reactions [6]. AVs are also capable of computing optimal decision-making solutions more swiftly than human drivers, thereby enhancing traffic efficiency [7]. Notably, AVs can enhance the capacity of roundabouts [8]. With the advent of AVs, whose amount is expected to exceed 50 million by 2024 [9], modern roundabout designs are increasingly accommodating AVs to address the challenges of safe control and interaction with HDVs [10]. Therefore, it is essential to ensure the safe operation of AVs in high-traffic roundabouts, particularly when they interact with HDVs [11]. Control strategies for connected autonomous vehicles (CAVs) are meticulously crafted to prioritize safety and efficiency in interactive driving situations [12]. Contemporary roundabout designs increasingly accommodate the operational needs of AVs, optimizing traffic flow and improving safety [13].
Existing studies on the interaction between AVs, CAVs, and HDVs in roundabouts can be broadly categorized into model-based and learning-based approaches. Model-based methods, such as game theory, have been widely used to model the decision-making process in interactive driving scenarios. Different driving behaviors in game processes can be modeled by adjusting the ratios of safety, efficiency, and driving comfort [14]. However, works using game process often rely on simplified environment and may not be able to handle the complex and dynamic nature of real-world roundabout scenarios [15, 16]. Other model-based frameworks face limitations due to the simplistic design of roundabouts, the low number of vehicles involved, and the limited generalization as only an abnormal roundabout is focused [17, 18].
Learning-based approaches, such as machine learning and deep reinforcement learning (DRL), have shown great potential in complex interactive driving in roundabouts. Machine learning techniques have been used to explore safe and efficient interaction between AVs and HDVs [19]. However, these machine learning-based methods often require a large amount of labeled data and struggle to generalize to unseen scenarios. DRL, on the other hand, allows for extensive exploration of strategies in complex interactive environments. Therefore, DRL is well-suited for the interactive driving in roundabouts. By learning from experience, DRL-based controllers can balance safety and efficiency in dense traffic scenarios.
Several DRL algorithms have been applied to autonomous driving, including Deep Deterministic Policy Gradient (DDPG)[20], Proximal Policy Optimization (PPO)[21], and deep Q-learning network (DQN)[22]. DDPG is well-suited for continuous action spaces, but its advantages may not be fully exploited in complex tasks with discrete decisions, such as driving in the roundabout[23]. PPO has been used to develop intelligent driving strategies that balance safety and efficiency in dense highway traffic [24]. However, PPO’s on-policy learning mode may be inefficient in utilizing historical data, which is crucial for approximating strategies in complex environments like roundabouts.
DQN has been effectively applied in various traffic simulations and is adept at tasks similar to roundabout navigation, such as intersection management [25]. DQN’s discrete action framework directly applies to tasks like proper lane selection in roundabouts without the need for action discretization. Additionally, DQN applies experience replay to relearn from past experiences, achieving more efficient use of historical data. Furthermore, DQN tends to be less computationally demanding [26]. The recently proposed Kolmogorov-Arnold Network (KAN) has improved performance compared to traditional multi-layer perceptrons [27]. KAN uses the B-spline function to replace the linear layer in the multi-layer perceptron, which can adaptively adjust the feature extraction method of each neuron. It can flexibly change the shape of the kernel function to learn key feature patterns related to the task, which improves the adaptability and generalization performance in different environments.
@@
To solve the complex driving in roundabouts, this paper proposes to integrate KAN with DQN (K-DQN) to enhance the decision-making and learning capabilities of AVs in complex roundabout scenarios. The K-DQN leverages the advantages of both DQN and KAN, enabling AVs to learn robust and efficient driving strategies through interaction with the environment. For conflict-free driving, we introduce an action inspector applied to time to collision (TTC) [28] to assess the relative collision risks between the AV and other HDVs. By replacing dangerous actions that may cause collisions with safe actions, our proposed method can decrease the ego vehicle collision rates with neighboring vehicles (NVs) during training. For proper lane selection, we introduce a route planner that considers the number of HDVs and the available free-driving space in each lane. For precise control of planned trajectories, we implement model predictive control (MPC) to allow the AV to navigate with precision and robustness.
The main contributions of this paper are summarized as follows:
  • We introduce a novel KAN-based DQN (K-DQN) to improve AVs’ decision-making in complex roundabouts. By integrating KAN with spline-based activation functions, the K-DQN boosts observation and decision-making capabilities, leading to enhanced training convergence, reduced collisions, and increased average speeds.
  • We propose an action inspector and a route planner to reduce collisions and enhance driving efficiency. The action inspector mitigates collision risks with HDVs, and the route planner optimizes lane selection for safety and space, significantly reducing collisions across various traffic flows compared to benchmarks.
@@
  • We integrate model predictive control (MPC) with K-DQN to convert planned actions into safe, smooth control commands, thereby enhancing journey safety and robustness. Our integrated solution adeptly manages diverse roundabout traffic flows, showing improved speed stability and efficiency over current benchmarks.
@@
The rest of this paper is organized as follows: Section II presents the problem statement and system structure; Section III describes the enhanced KAN-based DQN; Section IV introduces the action inspector and route planner; Section V presents the MPC design; Section VI provides the simulation results with analysis; Section VII draws the conclusion.
@@ II-A Problem Statement
The previous section underscored the significant attention given to studies on the interactive driving of AVs with HDVs in roundabouts. However, it is challenging in the integrated process of decision-making, path planning, and control for AVs to navigate roundabouts amid the varied complexity presented by HDVs. Interactions with HDVs, which can be randomly and densely distributed along both the inner and outer boundaries of roundabouts, frequently result in unexpected outcomes, such as conflicts and inefficient driving. Unlike straight or other curvy roads, roundabouts present unique challenges in making safe and efficient decisions due to their complex network of entrances and outlets. As illustrated in Fig. 1, there are four ports, each divided by a center white line into two half-ports—an entrance and an outlet. In this study, the right half-port of each is considered an entrance, while the left half-port serves as an outlet. The unpredictability of HDVs’ intended outlets and their various maneuvers exacerbates the challenges for AVs in ensuring safe passage. Moreover, satisfactory driving considerations encompass not only safety but also efficiency, which involves minimizing the total driving time of an AV to a designated outlet. The diverse and unpredictable actions of HDVs within the roundabout complicate the ability of AVs to achieve these goals simultaneously. This paper considers a double-lane roundabout without signal lights, featuring both HDVs and AVs as participants. In the modeled roundabout, HDVs are categorized into two types: those initially present within the roundabout and those merging into it. Three potential collision situations have been identified in Fig. 1, Fig. 2(a), and Fig. 2(b), respectively.
In Fig. 1, the first potential collision scenario is illustrated, where an AV is positioned at an entrance, while an HDV in the outer lane approaches the same entrance. This scenario leaves the AV with a decision influenced by the uncertain speed of the HDV. The AV has two options: to wait for the HDV to pass, thereby increasing its travel time, or to merge immediately, risking a collision, particularly if the relative velocity is low and the distance between the AV and HDV is minimal.
Figure 2 demonstrates the second potential collision scenario, in which the AV is navigating toward the same entrance from which an HDV is attempting to merge into the roundabout. This situation sets the stage for a possible collision if both vehicles attempt to occupy the entrance simultaneously.
Figure 2(b) illustrates the third potential collision scenario, which occurs when the AV wishes to exit from the inner lane but encounters an HDV in the adjacent outer lane. The AV has three options: overtaking the HDV with potential collision risks, following the HDV to reduce these risks, or staying in the current lane to wait for another opportunity to change lanes. These scenarios illustrate the high complexity and uncertainty that AVs face when engaging in interactive driving with surrounding HDVs in roundabouts.
Figure 1: The four-entrance, four outlet, and two-lane roundabout and an example of the first collision scenario with uncertainties of a autonomous vehicle.
Figure 2: Several challenging scenarios in the roundabout: (a) The second potential collision scenario; (b) The third potential collision scenario.
II-B System Structure
This paper primarily addresses the safety and efficiency of AVs from their entry to exit in roundabouts, taking into account the variable distribution and number of HDVs present. We consider varying traffic flows in the roundabout, from low to high density, and evaluate the system’s adaptability to these changes. To adapt to varying traffic flows, the proposed system requires a combination of techniques. First, the agent is required to adapt to various driving scenarios with different traffic densities. Second, the real-time changes in the environment is required to make safe and efficient decisions even in highly dynamic scenarios. Finally, the smooth and safe navigation is needed through the driving in roundabout.
To solve the above problems, the proposed KAN-based, conflict-avoidance, and proper-lane-detection DRL system is designed, as depicted in Fig. 3. This system comprises four key components: the environment, decision network, safety and efficiency mechanism, and robust control. The environment module updates the state of AVs based on the generated control commands and also generates a reward at each step to assess the value of the selected action. The decision network makes safe and efficient decisions for each step. The safety and efficiency mechanism comprises two components: the route planner and the action inspector. The former is activated during the initial merge into the roundabout, guiding AVs to select a lane that ensures sufficient driving space and safety. The latter assesses and mitigates collision risks with HDVs. For robust control, MPC is implemented to translate planned actions into safe and smooth control commands, thus enhancing safety and robustness throughout the journey.
Besides the normal driving, the system uses action inspector to detect their presence and take related response to emergency situations. The AV will choose actions that allow the emergency vehicle to pass safely and quickly, such as moving to the outer lane or slowing down. The route planner module will also adjust its path to minimize interference with the emergency vehicle’s trajectory.
III KAN-enhanced DQN Method
The K-DQN network includes a replay memory, a KAN-Q-network, and a target Q-network. The KAN-Q-network, featuring a robust and precise learning process and a Q-network, processes data from the environment to generate gradient-based updates. This network is tasked with observing the surrounding HDVs and calculating the Q-value for making safe and efficient decisions. The target Q-network has the same structure as the KAN-Q-network, but its parameters are updated less frequently. The target Q-network helps to mitigate the issue of oscillations or divergence in the learning process. The parameters of the target Q-network are periodically synchronized with those of the KAN-Q-network to ensure that the target values remain consistent with the learned Q-values. To provide a clear understanding of the variables used in the mathematical derivations of K-DQN, we summarize their descriptions in Table I. This table includes the variables related to the KAN-Q-network and the approximation theory used.
TABLE I: Variables used in the mathematical derivations
@@ III KAN-enhanced DQN Method
$s_{t}$ State at time step $t$
$a_{t}$ Action at time step $t$
$r_{t}$ Reward at time step $t$
@@
$\theta$ Parameters of the current Q-network
$\theta^{\prime}$ Parameters of the target Q-network
@@
$W$ Weights of the output layer
$b$ Bias of the output layer
@@
$C$ Constant in the approximation theory
$G$ Grid size in the approximation theory
$L_{b}$ Lipschitz constants for activation function
$L_{\text{spline}}$ Lipschitz constants for spline functions
@@ III-A Basic DQN
DQN combines deep learning with Q-learning, a reinforcement learning algorithm to address problems with high-dimensional state spaces. The core idea behind DQN is using a deep neural network to approximate the optimal action-value function $Q^{*}(s,a)$, which represents the maximum expected return for taking action $a$ in state $s$.
The two key components of DQN are experience replay and target networks. Experience replay stores the agent’s past experiences $(s_{t},a_{t},r_{t},s_{t+1})$ in a replay buffer, allowing the agent to learn from past experiences multiple times, which stabilizes the learning process. The target network is a separate neural network used to compute the target Q-values during training. The parameters of the target network are periodically synchronized with the main Q-network. This helps to stabilize the learning process by reducing the correlation between the target Q-values and the current Q-values.
In Q-learning, the Q-value update rule is given by
@@
where $Q(s_{t},a_{t})$ is the Q-value of executing action $a_{t}$ in state $s_{t}$ at time step $t$, $\alpha$ is the learning rate, $r_{t}$ is the immediate reward obtained after executing $a_{t}$ at time step $t$, $\gamma$ is the discount factor to balance the importance of immediate and future rewards, and $\max_{a}Q(s_{t+1},a)$ is the maximum Q-value of all possible actions in the next state $s_{t+1}$ at time step $t+1$.
During the training process, we define the target value $y$ as
@@
where $y$ is the ideal (target) Q-value of the current action calculated from the Bellman equation at time step $t$, and $\max_{a^{\prime}}Q(s_{t+1},a^{\prime};\theta^{\prime})$ is the maximum Q-value of all possible actions $a^{\prime}$ in the next state $s_{t+1}$ at time step $t+1$, estimated using parameters $\theta^{\prime}$. It is important to note that variables with subscript $t$ represent values at time step $t$, while variables without subscript $t$ represent values in a general sense.
We use the difference between the predicted Q-value $Q(s_{t},a_{t};\theta)$ at time step $t$ and the target value $y$ to design the loss function
@@
where $\theta$ is the parameters of the target network and $\mathbb{E}$ denotes the average expectation over all possible state-action pairs $(s_{t},a_{t})$ encountered during training.
In the DQN framework, the objective is to minimize the loss function $\mathcal{L}(\theta)$. Once the loss is calculated, the Q-network’s weights are updated in the opposite direction of the gradient to minimize the loss. The gradient $\frac{\partial\mathcal{L}}{\partial\theta}$ for updating the parameters in the DQN is derived as
$\frac{\partial\mathcal{L}}{\partial\theta}=\mathbb{E}\left[2(Q(s_{t},a_{t};\theta)-y)\frac{\partial Q}{\partial\theta}\right]$ (4)
with $y$ given in (2).
The gradients play a crucial role in updating the parameters of the DQN. They represent the rate of change of the loss function with respect to the Q-network’s parameters, indicating the direction in the parameter space that would lead to the steepest decrease in the loss function. By computing these gradients, we can determine how to adjust the parameters to minimize the loss function most effectively. Specifically, we update the parameters $\theta$ using the gradient descent algorithm, moving in the negative gradient direction:
@@
where $\alpha$ is the learning rate that controls the step size of each update. This update process is iterated until the loss function converges to a sufficiently low value, at which point the Q-network’s predictions will be close to the target values.

### v2

@@ Abstract
Safety and efficiency are crucial for autonomous driving in roundabouts, especially mixed traffic with both autonomous vehicles (AVs) and human-driven vehicles. This paper presents a learning-based algorithm that promotes safe and efficient driving across varying roundabout traffic conditions. A deep Q-learning network is used to learn optimal strategies in complex multi-vehicle roundabout scenarios, while a Kolmogorov-Arnold Network (KAN) improves the AVs’ environmental understanding. To further enhance safety, an action inspector filters unsafe actions, and a route planner optimizes driving efficiency. Moreover, model predictive control ensures stability and precision in execution. Experimental results demonstrate that the proposed system consistently outperforms state-of-the-art methods, achieving fewer collisions, reduced travel time, and stable training with smooth reward convergence.
@@ I Introduction
Roundabout designs vary by city scale [1], but typically feature a central island that vehicles must circulate around—clockwise or counterclockwise—facilitating smoother traffic flow and reducing interaction complexity [2]. As urban roadways evolve, roundabouts have improved traffic distribution and increased road capacity [3]. While they generally present fewer conflicts than other intersections [4], safety concerns intensify in high-traffic conditions due to a greater crash risk [5], especially in mixed traffic with both human-driven vehicles (HDVs) and autonomous vehicles (AVs).
Understanding human driving behavior in roundabouts, especially during entering, circulating, and exiting, is a key research focus. Merging and exiting require AVs to interact with surrounding HDVs, interpreting their intentions to make optimized decisions. Lane changes demand careful monitoring of HDVs and precise control to follow planned trajectories. Therefore, AVs navigating roundabouts must select lanes appropriately, monitor their environment, avoid collisions, and maintain precise control.
Research on autonomous driving in roundabouts has advanced from basic navigation to complex mixed-traffic scenarios. AVs reduce safety incidents caused by human errors like fatigue and distraction [6] and can make faster, optimal decisions [7]. They also enhance roundabout capacity [8]. With AVs expected to exceed 50 million on the road by 2030 [9], modern roundabouts are increasingly designed to support safe AV-HDV interactions [10, 11]. Control strategies for connected autonomous vehicles prioritize safety and efficiency [12], while current roundabout designs improve traffic flow and safety [13].
Control methods for autonomous driving in roundabouts have gained significant attention, with Model Predictive Control (MPC) and game theory being prominent model-based approaches. Game theory models decision-making by balancing safety, efficiency, and comfort [14], but often relies on simplified environments and struggles with real-world complexity [15]. Other model-based frameworks are limited by simplistic roundabout designs, few vehicles, and focus on abnormal cases [16]. While MPC effectively handles vehicle dynamics and safety constraints, current approaches still face challenges in complex real-world roundabouts [17, 18].
Learning-based methods, including machine learning and deep reinforcement learning (DRL), show strong potential for complex roundabout driving. Machine learning has been applied to AV-HDV interactions [19], but often requires extensive labeled data and struggles with generalization. DRL enables exploration of strategies in complex environments [20] and balances safety and efficiency in dense traffic [21]. Popular DRL algorithms include Deep Deterministic Policy Gradient (DDPG)[22], Proximal Policy Optimization (PPO)[23], and deep Q-learning (DQN)[24]. DDPG suits continuous actions but is less effective for discrete decisions like roundabout driving[25]. PPO achieves safe, efficient strategies in dense traffic [13] but its on-policy learning limits use of historical data, which is vital in complex environments like roundabouts [26] DQN has been effectively applied to traffic simulations, excelling in tasks like intersection management relevant to roundabout navigation [27]. Its discrete action framework suits lane selection without action discretization, and experience replay enhances learning efficiency using past data. Additionally, DQN is computationally efficient [28]. The recently proposed Kolmogorov-Arnold Network (KAN) outperforms traditional multi-layer perceptrons by replacing linear layers with adaptive B-spline functions, enabling flexible feature extraction and improved generalization across diverse environments [29, 30].
@@
To solve the complex driving in roundabouts, this paper proposes to integrate KAN with DQN (K-DQN) to enhance the decision-making and learning capabilities of AVs in complex roundabout scenarios [31]. The K-DQN leverages the advantages of both DQN and KAN, enabling AVs to learn robust and efficient driving strategies through interaction with the environment. For conflict-free driving, we introduce an action inspector applied to time to collision (TTC) [32] to assess the relative collision risks between the AV and other HDVs. By replacing dangerous actions that may cause collisions with safe actions, our proposed method can decrease the ego vehicle collision rates with neighboring vehicles (NVs) during training. For proper lane selection, we introduce a route planner that considers the number of HDVs and the available free-driving space in each lane. For precise control of planned trajectories, we implement MPC to allow the AV [33] to navigate with precision and robustness.
The main contributions of this paper are as follows:
  • We propose a novel K-DQN to enhance AV decision-making in complex roundabouts. Compared to the traditional neural networks, the unique spline-based activation functions of KAN enable more precise environmental learning and decision-making, resulting in better training convergence, lower collision rates, and higher average speeds.
  • Unlike prior methods that treat safety and efficiency separately, we introduce an integrated approach combining an action inspector and route planner. By merging TTC-based safety checks with density-aware lane selection, our method significantly reduces collisions while improving driving efficiency across diverse traffic conditions, compared to benchmarks.
@@
  • We enhance traditional DRL by integrating MPC with K-DQN, translating planned actions into safe, smooth controls. Our integrated solution adeptly manages diverse roundabout traffic flows, showing improved speed stability and efficiency over current benchmarks.
@@
The rest of this paper is organized as follows: Section II presents the problem statement and system structure; Section III describes the enhanced K-DQN; Section IV introduces the action inspector and route planner; Section V presents the MPC design; Section VI provides the simulation results with analysis; Section VII draws the conclusion.
Figure 1: A four-entrance, four-outlet, two-lane roundabout with the first collision scenario involving AV uncertainty.
Figure 2: Roundabout potential collision scenarios (a) and (b).
@@ II Problem Statement and System Structure
The previous section highlighted the focus on AV-HDV interaction in roundabouts. However, integrating decision-making, path planning, and control remains challenging due to the complexity of roundabout scenarios. Unlike straight or other curvy roads, roundabouts present unique challenges in making safe and efficient decisions due to their complex network of entrances and outlets. HDVs can be randomly and densely distributed along both the inner and outer boundaries of roundabouts, frequently resulting in unexpected outcomes such as conflicts and inefficient driving. As shown in Fig. 1, the roundabout has four ports, each split into an entrance (right) and an outlet (left). HDVs’ unpredictable maneuvers and unknown destinations pose challenges for AVs to ensure both safety and efficiency, defined here as minimizing travel time to the outlet. This study considers a signal-free, double-lane roundabout with both AVs and two types of HDVs: those already inside and those merging in.
Three potential collision situations have been identified in Fig. 1, Fig. 2(a), and Fig. 2(b), respectively. In Fig. 1, an AV at an entrance must decide whether to wait for an approaching HDV in the outer lane—potentially increasing delay—or merge immediately, risking a collision due to limited distance and uncertain HDV speed. Fig. 2(a) presents the reverse case: the AV is approaching an entrance while an HDV attempts to merge, creating a conflict if both proceed simultaneously. In Fig. 2(b), the AV intends to exit from the inner lane but encounters an HDV in the adjacent outer lane. It must choose between overtaking with risk, following to reduce conflict, or delaying the lane change. These scenarios highlight the complexity and uncertainty AVs face in interactive decision-making with HDVs in roundabouts.
This paper focuses on ensuring the safety and efficiency of AVs navigating roundabouts under varying HDV densities and traffic flows. The proposed system, illustrated in Fig. 3, addresses this by integrating adaptive decision-making, safety assurance, and robust control. The system comprises four components: environment, decision network, safety-efficiency mechanism, and robust control. The environment updates AV states and computes rewards based on control commands. The decision network selects actions that balance safety and efficiency. The safety-efficiency mechanism includes a route planner, which assists in lane selection during merging, and an action inspector that filters unsafe actions, especially in interactions with HDVs. For control execution, MPC ensures that the chosen actions are translated into smooth and reliable commands. In emergency scenarios, the action inspector identifies approaching emergency vehicles and triggers appropriate responses, such as yielding or changing lanes. The route planner concurrently adjusts the AV’s trajectory to minimize interference, ensuring compliance with safety norms.
TABLE I: Variables and description
@@
$s_{t},a_{t},r_{t}$ State, action, and reward at time step $t$
@@
$\theta$, $\theta^{\prime}$ Parameters of the current Q-network and target Q-network
@@
$W$, $b$ Weight and bias of the output layer
@@
$C$, $G$ Constant and Grid size in the approximation theory
$L_{b}$, $L_{\text{spline}}$ Lipschitz constants for activation, and spline functions
@@
III KAN-enhanced DQN Method
The K-DQN network consists of a replay memory, a KAN-Q-network, and a target Q-network. The KAN-Q-network processes environmental data to compute Q-values for safe, efficient decision-making, leveraging robust and precise learning. The target Q-network shares the same structure but updates less frequently to reduce learning instability. Its parameters are periodically synchronized with the KAN-Q-network. Key variables used in the K-DQN’s mathematical derivations are summarized in Table I.
@@ III-A Basic DQN
DQN integrates deep learning with Q-learning to handle high-dimensional state spaces. It uses a neural network to approximate the optimal action-value function $Q^{*}(s,a)$—the maximum expected return for taking action $a$ in state $s$. DQN relies on experience replay and target networks. Experience replay stores transitions $(s_{t},a_{t},r_{t},s_{t+1})$ for repeated learning, enhancing stability. The target network, updated periodically from the main Q-network, computes target Q-values to reduce correlations and stabilize training.
Let $Q(s_{t},a_{t})$ denote the Q-value for taking action $a_{t}$ in state $s_{t}$ at time step $t$. In Q-learning, the update rule for the Q-value is given by
@@
where $\alpha$ is the learning rate, $r_{t}$ is the immediate reward, $\gamma$ is the discount factor, and $\max_{a}Q(s_{t+1},a)$ is the maximum Q-value over actions in the next state $s_{t+1}$.
Let $y$ be the ideal (target) Q-value of the current action calculated from the Bellman equation at time step $t$. During the training process, $y$ is computed by
@@
where $\max_{a^{\prime}}Q(s_{t+1},a^{\prime};\theta^{\prime})$ is the maximum Q-value over actions $a^{\prime}$ in the next state $s_{t+1}$, estimated with parameters $\theta^{\prime}$. Note that the subscript $t$ denotes values at time $t$, while unsubscripted variables are general.
The loss function is defined on the difference between $Q(s_{t},a_{t};\theta)$ and $y$ as follows:
@@
where $\theta$ is the target network parameter, and $\mathbb{E}$ is the expectation over all state-action pairs $(s_{t},a_{t})$ during training.
In the DQN framework, the goal is to minimize the loss function $\mathcal{L}(\theta)$. After computing the loss, the Q-network’s weights are updated by descending the gradient to reduce $\mathcal{L}$. The gradient $\frac{\partial\mathcal{L}}{\partial\theta}$ used for this update is given by
$\frac{\partial\mathcal{L}}{\partial\theta}=\mathbb{E}\left[2(Q(s_{t},a_{t};\theta)-y)\frac{\partial Q}{\partial\theta}\right].$ (4)
Gradients are essential for updating the DQN parameters, as they measure how the loss function changes with respect to the Q-network’s parameters and indicate the direction of steepest descent. By computing these gradients, we identify how to adjust the parameters to effectively minimize the loss. Specifically, we update the parameters $\theta$ using gradient descent by moving in the negative gradient direction:
@@
where $\alpha$ is the learning rate controlling the update step size. This process repeats until the loss converges, aligning the Q-network’s predictions with the targets.
Basic DQN faces challenges in complex environments like roundabouts: training instability from correlated samples and moving targets, the exploration-exploitation trade-off, sensitivity to hyperparameters, and overestimation bias causing suboptimal policies. To balance exploration and exploitation, we use a dynamic $\varepsilon$-greedy strategy, gradually reducing $\varepsilon$ from 0.9 to 0.1 for broad exploration and stable refinement. KAN’s adaptive spline activations enhance feature representation and reduce hyperparameter sensitivity, improving robustness. Overestimation bias is mitigated by KAN’s accurate Q-value approximation and the action inspector’s real-time safety checks, preventing unsafe actions from skewing policy updates.